OpenStudy (anonymous):

integrat of eˆx/root(eˆx+1)

OpenStudy (thomas5267):

This? \[ \int \frac{e^x}{\sqrt{e^x+1}}\,dx \]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

Let \(e^x+1 = u\)